# A Minimalist Optimizer Design for LLM Pretraining

### v1

@@ 2.2 Momentum in the last layer
First-order momentum ($m^{t}$ in (2.3)) has been shown to effectively reduce the variance toward the true gradients; see Liu et al. (2020); Cutkosky and Mehta (2020). Although second-order momentum is used in adaptive optimizers such as Adam and RMSprop, more recent optimizers such as Muon have demonstrated remarkable success without second-order momentum. It is worth noticing that momentum is the major factor that introduces memory overhead in optimizer states, therefore we only consider first-order momentum in our optimizer design. Based on a minimalist optimizer design principle, we plan to investigate whether first-order momentum can be selectively used among different layers, while still leading to performance improvements without significantly increasing memory overhead.
Layer-wise gradient variance. Existing works already show the effectiveness of momentum in reducing variances (Liu et al., 2020; Cutkosky and Mehta, 2020). Naturally, if the gradients for different layers have different variances, we could achieve both memory efficiency and variance reduction by applying momentum to the layers with large variances. In the next paragraphs, we identify the layer-wise gradient variances and determine the most important layer that provides the largest performance gain when incorporating momentum; we also justify the effectiveness of different momentum parameters for different layers using a convergence analysis theory in Theorem 2.1.
@@
where $t=1,2,...,T$ is the iteration counter; $l=1,...,L$ represents different layers, i.e. we assume different layers contain different momentum with different hyperparameters. We have the following theoretical result (see Appendix A.5 for the proof).
@@
Suppose $\ell(\theta)$ in (2.1) is lower bounded by $\ell^{*}$, $\gamma$-smooth (i.e. $\nabla\ell(\theta)$ is Lipschitz continuous with constant $\gamma$), also the stochastic gradient is unbiased $\mathbb{E}_{\xi_{t}}\nabla_{\theta_{l}}\ell(\theta^{t};\xi_{t})=\nabla_{\theta_{l}}\ell(\theta^{t})$ and with bounded variance:
$\mathbb{E}_{\xi_{t}}\|\nabla_{\theta_{l}}\ell(\theta^{t};\xi_{t})-\nabla_{\theta_{l}}\ell(\theta^{t})\|^{2}\leq\sigma_{l}^{2}$ (2.9)
for all $l=1,...,L$ and $t=0,...,T-1$. With appropriate choice of hyperparameters $\eta_{l}\geq\eta$ (see (A.20) and (A.5)) and $\beta_{l}\leq 1-\delta$ ($\delta$ is an absolute constant), we have the following convergence result for update (2.8):
@@
where $\Delta_{1}=\ell(\theta^{1})-\ell^{*}$.
@@
Theorem 2.1 suggests that by taking different $\beta_{l}$ for different layers $l=1,...,L$, the convergence could be improved. The second term of the right-hand side of (2.10) takes the form
@@
where $x$ represents the layer-wise momentum hyperparameter $\beta_{l}$. It is straightforward to verify that $f(x)$ is decreasing in $(0,1-\delta]$, therefore an optimal strategy is to pick $\beta_{l}=1-\delta$. However from a memory-efficient point of view, only $\beta_{l}=0$ saves the memory of momentums.
Now, if $\sigma_{l}$ (variance of the gradient of the $l$-th layer) is significantly higher than other layers, then taking $\beta_{l}$ higher than other layers will result in better convergence, which is consistent with our experimental findings. On the other hand, if $\sigma_{l}$ is close to) zero, taking $\beta_{l}=0$ will provide memory-efficiency without harming the convergence too much. In particular, if the variances $\sigma_{l}\approx 0$ for $l=1,...,L-1$, one could take $\beta_{l}=0$ for $l=1,...,L-1$ and only keep the momentum for the last layer without significantly damaging the convergence rate.
Experiments on different gradient normalization with last-layer-momentum. It is observed in Table 2 that singular-value and column-wise normalization obtains better performance than the other normalizations for pretraining. We conduct another simple experiment to check the performance of the two types of normalization with last-layer-momentum. The results are summarized in Table 3, from which we observe that as the model size gets larger, both signular-value and column-wise normalizations + last-layer-momentum are matching Muon’s performance, and we choose column-wise due to the computational time concern in view of Table 4.
@@
Muon 28.86 22.20 16.70
Singular-val + last-layer-momentum 31.20 22.33 16.67
Column-wise + last-layer-momentum (ours) 30.81 22.57 16.32
Table 3: Evaluation perplexity of two normalizations (singular-value and column-wise) when combined with last-layer-momentum.
3 Proposed Algorithm
We now introduce our proposed optimizer, SCALE, detailed in Algorithm 1. The design of SCALE is motivated by empirical insights from our preceding experiments, which highlighted the importance of stabilizing updates in the last layer and controlling gradient scale via column-wise normalization. Accordingly, SCALE integrates two components: column-wise normalization of gradients and momentum updates restricted to the last layer. Despite its simplicity, SCALE is highly effective and requires only minimal modifications to standard implementations of Adam—typically just a few additional lines of code. As we will demonstrate, this lightweight design allows SCALE to outperform state-of-the-art baselines such as Adam and achieve performance competitive with Stable-SPAM and Muon, while using substantially less memory.
@@ 3 Proposed Algorithm
           $m_{l}^{t}=\beta_{t}m_{l}^{t-1}+(1-\beta_{t})\nabla_{\theta_{l}}\ell(\theta^{t};\xi_{t})$;
@@
           $m_{l}^{t}=\nabla_{\theta_{l}}\ell(\theta^{t};\xi_{t})$ (record the gradient directly);
@@
Connection to existing works. We now give a discussion on the connection of the proposed method to existing works. The proposed algorithm utilizes column-wise normalization, which is also discussed in Pethick et al. (2025). In particular, Pethick et al. (2025) uses momentum for all the layers, also singular-value normalization for most of layers except for the last layer, where they also propose an option of column-wise normalization (since Pethick et al. (2025) denotes the weight matrices as $W\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, their row-wise normalization is the same operation as our column-wise normalization); Our method also differs from SWAN (Ma et al., 2024) in the following aspects: first, SWAN utilizes both the row-wise and singular-value normalization while our method only utilizes column-wise normalization, indicating certain redundancy in existing works that apply multiple normalizations; second, SWAN applies Adam for the embedding and LM head layers (the first and last layers) which increases the memory overhead, whereas our approach only introduces first-order momentum for the last layer. We summarize the techniques used in different papers in Table 4.
Convergence of Algorithm 1. We remark here that the convergence provided in Pethick et al. (2025, Section 5) could be adopted to provide a convergence analysis for Algorithm 1, since Algorithm 1 differs from the general framework of Pethick et al. (2025, Algorithm 2) only by the fact that Algorithm 1 applies different momentum for different layers. However, directly applying such kinds of analysis will not be able to explain why using either normalization or momentum is advantageous as compared to vanila SGD, we choose not to present such a proof.
@@
[…]
@@
: Adafactor only records the 2nd order EMA of the column and row norms, effectively reducing the memory from $d_{\text{in}}\times d_{\text{out}}$ to $d_{\text{in}}+d_{\text{out}}$
: SWAN applies Adam for the first and last layer, following the practice of GaLore (see Ma et al. (2024, Appendix J)).
@@ 4 Experiments
In this section, we test the proposed SCALE algorithm for LLM pretraining. We test on pretraining LLaMA (60M, 130M, 350M, 1B, 7B) models on the C4 (Colossal Clean Crawled Corpus) dataset (Raffel et al., 2020).
Baselines. We compare the proposed SCALE with Adam, Stable-SPAM, Muon, GaLore, Fira, APOLLO(-Mini) and SWAN (see Section 1.1 for a more detailed introduction of these works). Stable-SPAM and Muon provide the better pretraining perplexities than Adam (Huang et al., 2025a; Jordan et al., 2024; Glentis et al., 2025). GaLore, Fira and APOLLO(-Mini) are memory-efficient optimizers that project the gradients to low rank. SWAN is an optimizer that uses both singular-value and row-wise normalized SGD except for the first and last layers, and it can achieve state-of-the-art [^6] We copy SWAN’s result directly from Ma et al. (2024) since we cannot replicate the result exactly due to code unavailability.. For Adam, Stable-SPAM, GaLore and Fira, we record their performance from Glentis et al. (2025), where a comprehensive benchmark is conducted for these methods.
It is worth noticing that GaLore, Fira, APOLLO(-Mini) and SWAN run Adam for the first and last layers for stable training. For the 60M model the first and last layers contain over 50% of the total network parameters, and around 40% for 130M model. For the 350M this goes down to less than 20% and for the 1B to about 10%. Therefore, for smaller models these methods have limited memory savings as compared to Adam, which trains a significant percentage of the network parameters.
Hyperparameters. For all LLaMA experiments, we follow Zhao et al. (2024) and set the sequence length to 256 and the batch size to 512, train using BF16 format and report evaluation perplexity as our metric. We also use a cosine learning rate with linear warm-up for the first 10% of the iterations. For low-rank optimizers (GaLore, Fira and APOLLO) we follow their suggested hyperparameters (including the rank) but tune the learning rates. For Muon we follow the implementation from Liu et al. (2025). For all the implemented methods, we conduct a hyperparameter search via wandb sweeps (see Appendix A.2 for details).
Results for 60M–1B models. We report the results of the evaluation perplexity in Table 5. We can see that the proposed algorithm outperforms existing memory efficient optimizers, also (nearly) matches the performance of Stable-SPAM and Muon, especially for larger models, with only 35–65% of the memory. We also contrast the performance with the memory consumed in Figure 1, where we can see that the proposed method is indeed at the Pareto frontier for optimal memory usage while maintaining the (near) state-of-the-art performance. This makes it a strong candidate for large-scale pretraining under memory constraints [^7] Notice that we construct our optimizer step by step via a minimalist design, and we already conduct the ablations throughout our journey (see Table 2 and 3, also Figure 3). We do not conduct further ablation studies in this section, but we include experiments on the learning rate sensitivity in Appendix A.4.. It also makes our method a minimalist baseline for benchmarking more sophisticated optimizers.
@@
Tokens 1.4B 2.6B 7.8B 13.1B
Adam 30.05 (0.35G) 23.13 (0.81G) 18.77 (2.21G) 16.52 (8.04G)
Stable-SPAM 28.77 (0.35G) 22.20 (0.81G) 16.80 (2.21G) 13.97 (8.04G)
Muon 28.86 (0.23G) 22.20 (0.54G) 16.70 (1.47G) 14.18 (5.36G)
GaLore 34.58 (0.28G) 25.31 (0.61G) 19.37 (1.59G) 15.57 (4.76G)
Fira 30.34 (0.28G) 22.96 (0.61G) 16.82 (1.59G) 15.10 (4.76G)
SWAN 30.00 (0.25G) 22.83 (0.46G) 17.14 (1.00G) 14.42 (3.20G)
APOLLO 30.94 (0.28G) 22.93 (0.61G) 16.75 (1.59G) 14.20 (4.76G)
APOLLO-Mini 31.85 (0.25G) 23.63 (0.46G) 17.11 (1.00G) 14.13 (3.20G)
SCALE (ours) 30.81 (0.15G) 22.57 (0.32G) 16.32 (0.80G) 14.25 (2.81G)
Table 5: Experiment results for pretraining LLaMA models on C4 dataset. The result marked is from Ma et al. (2024). The results marked are from Glentis et al. (2025).
Results for 7B model. We report the results of the evaluation perplexity in Table 6. Due to limited computational resources, we run a single experiment for SCALE with learning rate 1e-4 on 8$\times$×NVIDIA H200 141G GPUs. Following Zhu et al. (2025), we train for a total of 19.7B tokens, corresponding to 150K steps. We only compare with APOLLO(-Mini) which is the only baseline that reports the results for 7B model (with BF16 precision). From the table, we conclude that SCALE outperforms APOLLO(-Mini) in terms of both final evaluation perplexity and memory.
Steps 40K 80K 120K 150K
@@
Table 6: Experiment results for pretraining 7B LLaMA model on C4 dataset. The results marked are from Zhu et al. (2025).

### v2

@@ 2.2 Momentum in the last layer
First-order momentum ($m^{t}$ in (2.3)) has been shown to effectively reduce the variance toward the true gradients; see liu2020improved; cutkosky2020momentum. Although second-order momentum is used in adaptive optimizers such as Adam and RMSprop, more recent optimizers such as Muon have demonstrated remarkable success without second-order momentum. It is worth noticing that momentum is the major factor that introduces memory overhead in optimizer states, therefore we only consider first-order momentum in our optimizer design. Based on a minimalist optimizer design principle, we plan to investigate whether first-order momentum can be selectively used among different layers, while still leading to performance improvements without significantly increasing memory overhead.
Layer-wise gradient variance. Existing works already show the effectiveness of momentum in reducing variances (liu2020improved; cutkosky2020momentum). Naturally, if the gradients for different layers have different variances, we could achieve both memory efficiency and variance reduction by applying momentum to the layers with large variances. In the next paragraphs, we identify the layer-wise gradient variances and determine the most important layer that provides the largest performance gain when incorporating momentum; we also justify the effectiveness of different momentum parameters for different layers using a convergence analysis theory in Theorem 2.1.
@@
where t=1,2,,Tt=1,2,...,T is the iteration counter; l=1,,Ll=1,...,L represents different layers, i.e. we assume different layers contain different momentum with different hyperparameters. We have the following theoretical result (see Appendix A.10 for the proof).
@@
Suppose $\ell(\theta)$ in (2.1) is lower bounded by $\ell^{*}$, $\gamma$-smooth (i.e. $\nabla\ell(\theta)$ is Lipschitz continuous with constant $\gamma$), also the stochastic gradient is unbiased $\mathbb{E}_{\xi_{t}}\nabla_{\theta_{l}}\ell(\theta^{t};\xi_{t})=\nabla_{\theta_{l}}\ell(\theta^{t})$ and with bounded variance: $\mathbb{E}_{\xi_{t}}\|\nabla_{\theta_{l}}\ell(\theta^{t};\xi_{t})-\nabla_{\theta_{l}}\ell(\theta^{t})\|^{2}\leq\sigma_{l}^{2}$ for all l=1,,Ll=1,...,L and t=0,,T1t=0,...,T-1. With appropriate choice of hyperparameters $\eta_{l}\geq\eta$ (see (A.20) and (A.10)) and $\beta_{l}\leq 1-\delta$ ($\delta$ is an absolute constant), we have the following convergence result for update (2.7):
@@
where $\nabla_{l}^{t}=\nabla_{\theta_{l}}\ell(\theta^{t};\xi_{t})$ and $\Delta_{1}=\ell(\theta^{1})$-$\ell^{*}$.
@@
The left-hand side of (2.8) is the expectation of the stochastic gradient, measuring the stationarity. The right-hand side suggests an overall $\mathcal{O}(1/\sqrt{T})$ rate of convergence in terms of the iteration number $T$, meanwhile the summation indicates the error to each of the layer is aggregated, whose magnitude depends on the variance level $\sigma_{l}$ of the specific layer.
Theorem 2.1 suggests that by taking different $\beta_{l}$ for different layers l=1,,Ll=1,...,L, the convergence could be improved. The second term of the right-hand side of (2.8) takes the form
@@
where $x$ represents the layer-wise momentum hyperparameter $\beta_{l}$. It is straightforward to verify that $f(x)$ is decreasing in $(0,1-\delta]$, therefore an optimal strategy is to pick $\beta_{l}=1-\delta$. However from a memory-efficient point of view, only $\beta_{l}=0$ saves the memory of momentum.
Now, if $\sigma_{l}$ (variance of the gradient of the $l$-th layer) is significantly higher than other layers, then taking $\beta_{l}$ higher than other layers will result in better convergence, which is consistent with our experimental findings. On the other hand, if $\sigma_{l}$ is close to) zero, taking $\beta_{l}=0$ will provide memory-efficiency without harming the convergence too much. In particular, if the variances $\sigma_{l}\approx 0$ for l=1,,L1l=1,...,L-1, one could take $\beta_{l}=0$ for l=1,,L1l=1,...,L-1 and only keep the momentum for the last layer without significantly damaging the convergence rate.
Proof Sketch for Theorem 2.1. The proof follows liu2020improved, with the major difference being that liu2020improved proved the case with one single layer ($L=1$) and we extend it to multi-layer setting ($L>1$) where each layer has difference variance level.
First, define the auxiliary sequence by
$z_{l}^{t}:=\begin{cases}\theta_{l}^{t}&t=1\\ \frac{1}{1-\beta_{l}}\theta_{l}^{t}-\frac{\beta_{l}}{1-\beta_{l}}\theta_{l}^{t-1}&t\geq 2\end{cases}$ (2.9)
we can verify two important properties: $z_{l}^{t+1}-z_{l}^{t}=-\eta_{l}g_{l}^{t}$ and $z_{l}^{t}-\theta_{l}^{t}=-\frac{\beta_{l}}{1-\beta_{l}}\eta_{l}m_{l}^{t-1}$ (Lemma A.3). Using $\gamma$-Lipschitz smooth, we can expand the function value difference via
$\mathbb{E}_{\xi_{t}}\left[\ell\left(z^{t+1}\right)\right]\leq\ell\left(z^{t}\right)+\underbrace{\mathbb{E}_{\xi_{t}}\left[\left\langle\nabla\ell\left(z^{t}\right),z^{t+1}-z^{t}\right\rangle\right]}_{(a)}+\frac{\gamma}{2}\underbrace{\mathbb{E}_{\xi_{t}}\left[\left\|z^{t+1}-z^{t}\right\|^{2}\right]}_{(b)}.$
Expanding the term (a) on the right-hand side using the update using the above properties of $z_{l}^{t}$, we will arrive at a $\mathcal{O}(\sum_{t=1}^{T}\eta_{t}\sum_{l=1}^{L}\mathbb{E}\|\nabla_{l}^{t}\|^{2})$ term (which is the left-hand side of (2.8)), and some addition term. The additional term, combining with (b), will be bounded via Lemma A.1 and A.2, and together form the right hand side of (2.8).
@@
Adam 30.05 23.13 18.77
Adam (Stable-SPAM) 28.77 22.20 16.80
Singular-val (NS) + mmt-last 31.20 22.33 16.67
Column-wise + mmt-last (ours) 30.81 22.57 16.32
Table 3: Evaluation perplexity of two normalizations (singular-value and column-wise) when combined with last-layer-momentum (mmt-last).
Experiments on different gradient normalization with last-layer-momentum. It is observed in Table 2 that singular-value and column-wise normalization obtains better performance than the other normalizations for pretraining. We conduct another experiment to check the performance of the two types of normalization with last-layer-momentum. The results are summarized in Table 3, where we observe that as the model size gets larger, both singular-value and column-wise normalizations + last-layer-momentum are matching Adam’s performance, and we choose column-wise due to the computational time concern in Table 3.
3 The SCALE Algorithm
The above investigation leads to the following simple optimizer, SCALE, detailed in Algorithm 1. The design of SCALE is motivated by empirical insights from our preceding experiments, which highlighted the importance of stabilizing updates in the last layer and controlling gradient scale via column-wise normalization. Accordingly, SCALE integrates two components: column-wise normalization of gradients and momentum updates restricted to the last layer. Despite its simplicity, SCALE is highly effective and requires only minimal modifications to standard implementations of Adam—typically just a few additional lines of code. As we will demonstrate, this lightweight design allows SCALE to outperform state-of-the-art baselines such as Adam and achieve performance competitive with Stable-SPAM and Muon, while using substantially less memory.
@@ 3 The SCALE Algorithm
    $m_{l}^{t}=\beta_{t}m_{l}^{t-1}+(1-\beta_{t})g_{l}^{t}$;
@@
    $m_{l}^{t}=g_{l}^{t}$ (record the gradient directly);
@@
Connection to existing works. We now give a discussion on the connection of the proposed method to existing works. The proposed algorithm utilizes column-wise normalization, which is also discussed in pethick2025scion. pethick2025scion uses momentum for all the layers and applies column-wise normalization as an alternative to singular-value normalization only for the last layer. Our method also differs from SWAN (ma2024swan) in the following aspects: first, SWAN utilizes both the row-wise and singular-value normalization while our method only utilizes column-wise normalization, indicating certain redundancy in existing works that apply multiple normalizations; second, SWAN applies Adam for the embedding and LM head layers (the first and last layers) which increases the memory overhead, whereas our approach only introduces first-order momentum for the last layer. We summarize the techniques used in different papers in Table 4.
Convergence of Algorithm 1. We remark here that the convergence provided in pethick2025scion could be adopted to provide a convergence analysis for Algorithm 1, since Algorithm 1 differs from the general framework of pethick2025scion only by the fact that Algorithm 1 applies different momentum for different layers. However, directly applying such kinds of analysis will not be able to explain why using either normalization or momentum is advantageous as compared to vanilla SGD, we choose not to present such a proof.
@@
SWAN First/Last Layer First/Last Layer 14.52
APOLLO Rank-256 Rank-256 16.14
APOLLO-Mini Rank-1 Rank-1 14.53
@@
: SWAN, APOLLO and APOLLO-Mini apply Adam for the first and last layer.
@@ 4 Experiments
In this section, we test the SCALE algorithm for LLM pretraining. We test on pretraining LLaMA (60M, 130M, 350M, 1B, 7B) models on the C4 (Colossal Clean Crawled Corpus) dataset (raffel2020exploring) and report evaluation perplexity as our metric. For all ours experiments, we follow the hyperparameter settings of zhao2024galore, see Appendix A.2 for more details.
Baselines. We compare with Adam and its stabilized version, noted as Adam (Stable-SPAM), which performs momentum resets and clips spiked gradients (huang2025stablespam). Among memory efficient optimizers, we compare with Muon, GaLore, Fira, APOLLO(-Mini) and SWAN (see Section 1.1 for a more detailed introduction of these works). GaLore, Fira and APOLLO(-Mini) compress the Adam states by projecting the gradients to low rank. SWAN is an optimizer that uses both singular-value and row-wise normalization, except for the first and last layers [^5] We copy SWAN’s result from ma2024swan since we cannot replicate them due to code unavailability..
It is worth noticing that GaLore, Fira, APOLLO(-Mini) and SWAN run Adam for the first and last layers for stable training. For the 60M model the first and last layers contain over 50% of the total network parameters, and around 40% for 130M model. For the 350M this goes down to about 20% and for the 1B to about 10%. Therefore, for smaller models these methods have limited memory savings as compared to Adam, which is used to optimize a significant percentage of the network parameters.
Results for 60M–1B models. We report the results of the evaluation perplexity in Table 5. We can see that, despite its simplicity, SCALE outperforms existing memory efficient optimizers, also (nearly) matches the performance of Adam (Stable-SPAM) and Muon, especially for larger models, with only 35–65% of the memory. We also contrast the performance with the memory consumed in Figure 1, where we can see that the proposed method is indeed at the Pareto frontier for optimal memory usage while maintaining the (near) state-of-the-art performance. This makes it a strong candidate for large-scale pretraining under memory constraints [^6] Notice that we construct our optimizer step by step via a minimalist design, and we already conduct the ablations throughout our journey (see Table 2 and 3, also Figure 4). We do not conduct further ablation studies in this section, but we include experiments on throughput analysis, additional architectures, overtraining, finetuning and learning rate sensitivity in the Appendix sections A.3, A.4, A.5, A.6 and A.7, respectively..
@@
Tokens 1.4B 2.6B 7.8B 20B
Adam 30.05 (0.35G) 23.13 (0.81G) 18.77 (2.21G) 15.79 (8.04G)
Adam (Stable-SPAM) 28.77 (0.35G) 22.20 (0.81G) 16.80 (2.21G) 13.30 (8.04G)
Muon 28.86 (0.23G) 22.20 (0.54G) 16.70 (1.47G) 13.67 (5.36G)
GaLore 34.58 (0.28G) 25.31 (0.61G) 19.37 (1.59G) 15.05 (4.76G)
Fira 30.34 (0.28G) 22.96 (0.61G) 16.82 (1.59G) 14.36 (4.76G)
SWAN 30.00 (0.25G) 22.83 (0.46G) 17.14 (1.00G) -
APOLLO 30.94 (0.28G) 22.93 (0.61G) 16.75 (1.59G) 14.28 (4.76G)
APOLLO-Mini 31.85 (0.25G) 23.63 (0.46G) 17.11 (1.00G) 13.48 (3.20G)
SCALE (ours) 30.81 (0.15G) 22.57 (0.32G) 16.32 (0.80G) 13.49 (2.81G)
Table 5: Experiment results for pretraining LLaMA models on C4 dataset. The result marked is from ma2024swan. The results marked for model sizes 60M-350M are from glentis2025scalable. All models are trained up-to the Chinchilla optimal number of tokens (hoffmann2022training). For Fira and APOLLO 1B runs we encountered training divergence with their default learning rates, results reported are with reduced learning rate. Among the memory-efficient optimizers we highlight the best-performing for each model size in terms of perplexity and memory. Note that SWAN does not provide a 1B model result with Chinchilla optimal tokens, but only for 13B tokens. In that setting we also achieve a superior perplexity of 14.25 (2.81G memory), compared to 14.42 (3.20G memory) of SWAN.
Steps 40K 80K 120K 150K (final)
@@
Table 6: Results for pretraining 7B LLaMA model on C4 dataset. From zhu2025apollo.
[…]